# Dynamic Gaussians Mesh: Consistent Mesh Reconstruction from Monocular Videos

### v2

@@ 4 Experiments
We introduce the dataset we evaluate on and our implementation details in Sec. 4.1 and Sec. 4.2. In Sec. 4.3, we evaluate our method compared with other baselines. Regarding the baselines, we primarily compare with the following works: D-NeRF [27], K-Plane [6], HexPlane [1], TiNeuVox [5]. Among these baselines, D-NeRF [27] and TiNeuVox [5] learn a deformation field to transform the dynamic scene into the canonical space, while other methods like HexPlane [1] and K-Plane [6] utilize volume factorization to learn a compact 4D feature volume. Note that all these baselines do not model scene geometry explicitly, to compare the mesh quality, we extract mesh from either an implicit density field or an occupancy field queried from the 3D Gaussians. In Sec. 4.4, we studied several regularization factors in our method and provided quantitative results of how each factor affects the final performance. We also show several applications such as ray-tracing and texture editing of the time-consistent mesh we export from DG-Mesh.
@@ 4.1 Datasets
The D-NeRF [27] synthetic dataset does not provide mesh ground truth information. Therefore, we rendered a synthetic dataset containing six dynamic scenes to compare the mesh reconstruction quality quantitatively. Each scene includes 200 frames of a moving object with the ground truth camera parameters and images, as well as the ground truth mesh under each time frame. Camera views are evenly distributed around the target objects on the sphere or the upper sphere. We also evaluate our method on D-NeRF synthetic dataset and provide the visualization and comparison with other baselines in Fig. 5. For real data, we run our method on the Nerfies dataset [23] which contains monocular videos of daily life objects captured using hand-held cameras. We provide the visualization of the reconstructed mesh as well as the mesh rendering images in Fig. 7. We also tested our results on the Unbiased4D dataset [9], which provides one sequence of real monocular video captures of a moving cactus toy.
Figure 5: We compare the mesh reconstruction and rendering results of our method and other baselines on the D-NeRF dataset. For each object, we visualize the results under two different time steps and different viewing angles. Our method outperforms others by producing more smooth and detailed structures.
@@ 4.2 Implementation Details
We used PyTorch [25] to implement our method and utilized the implementation from 3D Gaussian Splatting [13] for differential Gaussian rasterization. Our model was trained for a total of 50,000 iterations on a single RTX 3090Ti. To better initialize deformable 3D Gaussians, we first trained the canonical Gaussians for $3k$ iterations while keeping the forward and backward deformation network fixed. This helped to retain relatively stable positions and shapes of 3D Gaussians under the canonical space. After $5k$ iterations, we introduce the DPSR and differentiable Marching Cubes to extract the mesh geometry from the Gaussian points. We perform Gaussian-Mesh Anchoring every $100$ iteration during training. For the mesh rasterization, we use Nvdiffrast, to query the vertex color of the deformed mesh, we project the location of the deformed vertex back to the canonical space and query a time-dependent appearance module to obtain the color. We found optimizing the vertex color in canonical space produces better results compared with querying color under different deformed spaces. The final supervision comes from the rendering loss from both the splatted Gaussian images and the rendered mesh images.
Figure 6: We provide the result comparison of our method with other baselines in our rendered dataset. Our method gives better mesh rendering quality as well as geometry. For objects such as the flying bird, other methods fail to recover the thin geometry of the bird wings. Instead, our method produces satisfying geometry. Our method also produces better surface rendering.
@@ 4.3 Results and Comparisons
For the D-NeRF dataset, we present our numerical results and comparisons in Tab. 2 and mesh visualization in Fig. 5. We compare the mesh rendering PSNR, SSIM, and LPIPS scores for each object in the dataset. Since the D-NeRF dataset does not provide ground truth mesh, we do not numerically evaluate our performance of the mesh quality. In Fig. 5, we highlight the part where our method outperforms the other baselines and recovers the highly detailed structures. For our rendered dataset with mesh ground truth, we provide the numerical results and comparisons in Tab. 1 and visualizations in Fig. 6. Regarding evaluation metrics of mesh quality, we chose two different metrics to measure the accuracy of our mesh reconstruction: the Chamfer Distance (CD) and the Earth Mover Distance (EMD) to measure the displacement between the reconstructed mesh and the ground truth. As shown in Tab. 1, our method achieves lower CD and EMD scores compared with other baselines in general, indicating our method reconstructs the highest quality mesh. In Fig. 6, we show the visualization of our reconstructed mesh compared to other baselines in our dataset. In challenging areas, such as thin structures like the wings of the bird, our method still recovers high-quality geometry and surfaces, while other methods produce floaters or oversized meshes.
@@
Ours 30.400 0.968 0.055 30.208 0.974 0.051 28.951 0.959 0.065 25.460 0.959 0.084
Table 2: We measure the mesh rendering PSNR, SSIM and LPIPS score of all the baselines and our method on the D-NeRF dataset. Our method has the highest surface rendering quality compared with all the other baselines.
@@ 4.4 Ablation Study
Laplacian Smoothness Regularizer. We study the effect of laplacian regularizer and its different performance under different loss weights in this part. Laplacian regularization term calculates the average square Euclidean difference between the vertex and its one-ring neighbors, which discourages deformation that will introduce large surface curvature change. We measure the final mesh rendering quality (PSRN, SSIM, LPIPS) under different laplacian loss weights. As shown in Tab. 3, when $w_{lap}=1000$, the network produces the highest mesh rendering quality. With appropriate laplacian regularization, our method recovers a smoother surface.
@@
Gaussian-Mesh Anchoring. We study the effect of our proposed Gaussian-Mesh Anchoring procedure. Since the original densification and pruning mechanism of 3DGS tends to split more Gaussian in hard areas where the geometry and appearance are more complicated. The original Gaussian center distribution is not even on the object’s surface. As discussed in Sec. 3.3, DPSR assumes uniformly sampled points among the space. An uneven point cloud on the surface will produce the wrong distance field and harm the optimization process. Our Gaussian-Mesh Anchoring performs densification and pruning that will result in a more evenly distributed Gaussian point and rely on the backward deformation network to project it back to the canonical space. As shown in Fig. 2, after applying Gaussian-Mesh Anchoring, the Gaussian centers are more evenly distributed on the object’s surface. We provide quantitative results of how the Gaussian-Mesh Anchoring procedure allows the network to produce better geometry in Tab. 4. Without Gaussian-Mesh Anchoring, the reconstructed CD and EMD are both higher, and the mesh rendering is also not as good as it is with the anchoring.
Figure 7: We provide the reconstructed mesh visualization and mesh rendering image of our method on real-world datasets including the Nerfies dataset with monocular videos of daily life objects captured using hand-held cameras; the Unbiased4D dataset contains one sequence of real monocular video captures of a moving toy.
Figure 8: Two types of application can be done with time-consistent mesh: 1. Ray-tracing in the rendering engine. 2. Texture editing: With correspondence across time frames, we just need to edit the first frame in a sequence and the change will be automatically applied to the rest of the sequence.
CD $\downarrow$ EMD $\downarrow$ $\text{PSNR}_{m}$uparrow SSIM $\uparrow$ LPIPS $\downarrow$
w./o. anchor 0.685 0.146 31.283 0.921 0.048
w./ anchor 0.516 0.128 32.734 0.9180 0.041
Table 4: We turn the Gaussian-Mesh Anchoring on and off and measure the reconstructed mesh quality. The results indicate that the anchoring procedures improve the performance by guiding the Gaussian points to deform more evenly on the surface.
Applications DG-Mesh provides dense correspondence, which is particularly useful for a variety of downstream applications such as shape manipulation and texture editing. We showcase two examples of mesh editing applications enabled by our method in Fig. 8. We can insert the extracted mesh into new scene and perform ray-tracing. Given our time-consistent mesh under the first frame, we edit the mesh vertex’s color in the first frame by painting color to its surface; the same vertex color and pattern are propagated to the following frames. This editing process is straightforward to use.
@@ 5 Conclusion
We introduce Dynamic Gaussians Mesh (DG-Mesh), a framework to reconstruct high-fidelity mesh and perform motion tracking from monocular videos in dynamic scenes. Leveraging the recent advancements in 3D Gaussian Splatting, DG-Mesh learns a deformable 3D Gaussians set and builds a novel time-consistent mesh on top of it. DG-Mesh outperforms previous methods in reconstructing intricate structures like bird wings and horse legs, which are typically difficult to capture. Moreover, DG-Mesh’s ability to provide cross-frame correspondences enables various downstream applications, such as texture editing on dynamic objects, making it a versatile tool with great potential in graphics applications. We will release our code upon paper publication.

### v3

@@ 4 Experiments
We introduce the dataset we evaluate on and our implementation details in Section 4.1 and Section 4.2. In Section 4.3, we evaluate our method compared with other baselines. Regarding the baselines, we primarily compare with the following works: D-NeRF (Pumarola et al., 2021), K-Plane (Fridovich-Keil et al., 2023), HexPlane (Cao & Johnson, 2023), TiNeuVox (Fang et al., 2022), 4D-GS (Wu et al., 2023a), Deformable-GS (Yang et al., 2023d), 4DGS (Yang et al., 2023c), and SC-GS (Huang et al., 2024). Among these baselines, D-NeRF (Pumarola et al., 2021) and TiNeuVox (Fang et al., 2022) learn a deformation field to transform the dynamic scene into the canonical space, while other methods like HexPlane (Cao & Johnson, 2023) and K-Plane (Fridovich-Keil et al., 2023) utilize volume factorization to learn a compact 4D feature volume. 4D-GS (Wu et al., 2023a), Deformable-GS (Yang et al., 2023d) 4DGS (Yang et al., 2023c), and SC-GS (Huang et al., 2024) use 3D Gaussians as their representations and model their dynamics over time. Note that most of these baselines do not model scene geometry explicitly. To compare the mesh quality, we extract mesh from either an implicit density field or an occupancy field queried from the 3D Gaussians. In Section 4.4, we studied several regularization factors in our method and provided quantitative results of how each factor affects the final performance.
@@ 4.1 Datasets
We evaluate our method on the D-NeRF synthetic dataset and provide the visualization and comparison with other baselines in Figure 5. Since the D-NeRF (Pumarola et al., 2021) dataset does not provide mesh ground truth information, we rendered a synthetic dataset containing six dynamic scenes to compare the mesh reconstruction quality quantitatively. Each scene includes 200 frames of a moving object with the ground truth camera parameters and images, as well as the ground truth mesh under each time frame. Camera views are evenly distributed around the target objects on the sphere or the upper sphere. The reconstruction visualization is provided in Figure 6, and full evaluation visualizations can be found in the Appendix L.
For real data evaluation, we run our method on the Nerfies dataset (Park et al., 2021a) , the Dycheck’s dataset (Gao et al., 2022) and the Unbiased4D dataset (Johnson et al., 2023), all of which contains monocular videos of everyday deformable objects captured using handheld cameras. To demonstrate the adaptability of DG-Mesh to different dynamic capturing setups, we also test our method on the NeuralActor (Liu et al., 2021) dataset, which features multi-view videos of moving humans. Additionally, we captured several dynamic scenes using a smartphone.
@@ 4.2 Implementation Details
We utilize the implementation from 3D Gaussian Splatting (Kerbl et al., 2023) for differential Gaussian rasterization. Our model was trained for a total of 50,000 iterations on a single RTX 3090Ti. To better initialize deformable 3D Gaussians, we first trained the canonical Gaussians for $3k$ iterations while keeping the forward and backward deformation network fixed. This helped to retain relatively stable positions and shapes of 3D Gaussians under the canonical space. After $5k$ iterations, we introduce the DPSR and differentiable Marching Cubes to extract the mesh geometry from the Gaussian points. We perform Gaussian-Mesh Anchoring every $100$ iteration during training. For the mesh rasterization, we use Nvdiffrast. In order to query the vertex color of the deformed mesh, we project the location of the deformed vertex back to the canonical space and query a time-dependent appearance module to obtain the color. We found optimizing the vertex color in canonical space produces better results compared with querying color under different deformed spaces. The final supervision comes from the rendering loss from both the splatted Gaussian images and the rendered mesh images. More implementation details can be found in Appendix B.
@@ 4.3 Results and Comparisons
For the D-NeRF dataset, we present our numerical results and comparisons in Table 2 and mesh visualization in Figure 5. We compare the mesh rendering PSNR, SSIM, and LPIPS scores for each object in the dataset. Since the D-NeRF dataset does not provide ground truth mesh, we do not numerically evaluate our performance of the mesh quality. In Figure 5, we highlight the part where our method outperforms the other baselines and recovers the highly detailed structures. For the DG-Mesh dataset with mesh ground truth, we provide the numerical results and comparisons in Table 1 and visualizations in Figure 6. Regarding evaluation metrics of mesh quality, we chose two metrics to measure our mesh reconstruction’s accuracy: the Chamfer Distance (CD) and the Earth Mover Distance (EMD) to measure the displacement between the reconstructed mesh and the ground truth. As shown in Table 1, our method achieves lower CD and EMD scores compared with other baselines, indicating our method reconstructs the highest quality mesh. In Figure 6, we show the visualization of our reconstructed mesh compared to other baselines in our dataset. In challenging areas, such as thin structures like the wings of the bird, our method still recovers high-quality geometry and surfaces, while other methods produce floaters or oversized meshes. For real evaluation, we provide the quantitative results on the Nerfies dataset in Table 5 and qualitative results in Appendix G. Our method achieves the highest mesh query speed while delivering rendering quality comparable to the baseline. Performance on other monocular datasets is detailed in the supplementary materials, including the Unbiased4D dataset in Appendix J, the Dycheck dataset in Appendix H, and the self-captured iPhone dataset in Appendix I. Additionally, we present qualitative evaluations on the NeuralActor dataset in Appendix K, demonstrating the broader applicability of DG-Mesh to multiview setups.
@@
DG-Mesh 30.400 0.968 0.055 30.208 0.974 0.051 28.951 0.959 0.065 25.460 0.959 0.084
Table 2: We measure the mesh rendering PSNR, SSIM and LPIPS score of the baselines and our method on the D-NeRF dataset. Our method has the highest surface rendering quality compared with all the other baselines.
@@ 4.4 Ablation Study
Laplacian Smoothness Regularizer. We study the effect of laplacian regularizer and its different performance under different loss weights in this part. Laplacian regularization term calculates the average square Euclidean difference between the vertex and its one-ring neighbors, which discourages deformation that will introduce large surface curvature change. We measure the final mesh rendering quality (PSRN, SSIM, LPIPS) under different laplacian loss weights. As shown in Table 3, when $w_{lap}=1000$, the network produces the highest mesh rendering quality. With appropriate laplacian regularization, our method recovers a smoother surface.
@@
Gaussian-Mesh Anchoring. We study the effect of our proposed Gaussian-Mesh Anchoring procedure. Since the original densification and pruning mechanism of 3DGS tends to split more Gaussian in hard areas where the geometry and appearance are more complicated. The original Gaussian center distribution is not even on the object’s surface. As discussed in Section 3.3, DPSR assumes uniformly sampled points among the space. An uneven point cloud on the surface will produce the wrong distance field and harm the optimization process. Our Gaussian-Mesh Anchoring performs densification and pruning that will result in a more evenly distributed Gaussian point and rely on the backward deformation network to project it back to the canonical space. As shown in Figure 2, after applying Gaussian-Mesh Anchoring, the Gaussian centers are more evenly distributed on the object’s surface. We provide quantitative results of how the Gaussian-Mesh Anchoring procedure allows the network to produce better geometry in Table 5. Without Gaussian-Mesh Anchoring, the reconstructed CD and EMD are both higher.
Method CD $\downarrow$ EMD $\downarrow$ $\text{PSNR}_{m}$uparrow
w./o. anchor 0.685 0.146 31.283
w./ anchor 0.516 0.128 32.734
Table 4: Ablation study on the Gaussian-Mesh Anchoring. The anchoring allows mesh to be better aligned with the ground truth surface.
Method PSNR $\uparrow$ $\text{PSNR}_{m}$uparrow $\text{FPS}_{m}$uparrow
Deformable-GS 28.310 - 0.10
DG-Mesh - 27.238 10.9
Table 5: Real scene evaluation on the Nerfies dataset. Our method maintains high mesh querying FPS and comparable image rendering quality to the baseline.
Figure 7: We provide the reconstructed mesh visualization and mesh rendering image of our method on the Nerfies real-world datasets. Our method recovers high-quality and consistent mesh surfaces across multiple time frames, showcasing its robustness and effectiveness in handling dynamic scenes.
@@ 5 Conclusion
We introduce Dynamic Gaussians Mesh (DG-Mesh), a framework to reconstruct high-fidelity mesh and perform motion tracking from dynamic observations. Leveraging the recent advancements in 3D Gaussian Splatting, DG-Mesh learns a deformable 3D Gaussians set and builds a novel time-consistent mesh on top of it. DG-Mesh outperforms previous methods in reconstructing intricate structures like bird wings and horse legs, which are typically difficult to capture. Moreover, DG-Mesh’s ability to provide cross-frame correspondences enables various downstream applications, such as texture editing on dynamic objects, making it a versatile tool with great potential in graphics applications. We will release our code upon paper publication.